OpenStudy (anonymous):

Use completing the square to convert from standard form into vertex form: f(x)=x^2+8x=10

OpenStudy (thomas5267):

Note that \(\left(x+4\right)^2=x^2+8x+16\).

OpenStudy (anonymous):

still a bit confused what does that mean